# Information Density Bounds for Privacy

### v1

@@ Abstract
This paper explores the implications of guaranteeing privacy by imposing a lower bound on the information density between the private and the public data. We introduce an operationally meaningful privacy measure called pointwise maximal cost (PMC) and demonstrate that imposing an upper bound on PMC is equivalent to enforcing a lower bound on the information density. PMC quantifies the information leakage about a secret to adversaries who aim to minimize non-negative cost functions after observing the outcome of a privacy mechanism. When restricted to finite alphabets, PMC can equivalently be defined as the information leakage to adversaries aiming to minimize the probability of incorrectly guessing randomized functions of the secret. We study the properties of PMC and apply it to standard privacy mechanisms to demonstrate its practical relevance. Through a detailed examination, we connect PMC with other privacy measures that impose upper or lower bounds on the information density. Our results highlight that lower bounding the information density is a more stringent requirement than upper bounding it. Overall, our work significantly bridges the gaps in understanding the relationships between various privacy frameworks and provides insights for selecting a suitable framework for a given application.
Index Terms:
Privacy, information density, pointwise maximal leakage, pointwise maximal cost, local differential privacy, information privacy, gain function, cost function.
@@ I-A Motivations
There are two key motivations behind this paper. First, we seek to provide an in-depth understanding of the implications of defining privacy through a lower bound on the information density. To this end, we introduce an operationally meaningful privacy measure, which we call pointwise maximal cost (PMC). We demonstrate that imposing an upper bound on PMC is equivalent to imposing a lower bound on the information density. PMC may be thought of as a complementary definition to PML. More precisely, while PML considers opportunistic adversaries aiming to maximize gain functions, PMC focuses on risk-averse adversaries aiming to minimize arbitrary non-negative cost functions.
@@ I-B Contributions and Summary of the Main Results
Next, we argue that PMC can alternatively be defined as the multiplicative decrease in the expected value of an arbitrary non-negative cost function upon observing an outcome $y$ of the privacy mechanism. This setup is referred to as the cost function model of PMC. Essentially, we prove an equivalence between the randomized function model and the cost function model of PMC (Theorem 2). This equivalence signifies a parallelism with PML since a similar equivalence between the gain function model and the randomized function model of PML was previously established in [12, Thm. 2].
In Sections III-C and III-D, we explore the connections between PMC and other information measures designed with risk-averse adversaries in mind. These are maximal cost leakage [22, Def. 11], maximal realizable cost [22, Def. 12], and maximal guesswork leakage [23]. By carefully examining these definitions and their formulations, we observe that they are often more cumbersome and complex without strengthening the privacy model. For instance, both maximal cost leakage and maximal realizable cost involve a supremum over a Markov chain, which we demonstrate can be omitted. By contrast, we show that PMC offers a streamlined yet powerful and flexible formulation that effectively encompasses these alternative definitions. It is important to note that our discussions here are restricted to finite random variables, as the other measures have only been defined in this context.
@@
In Section V, we explore the interactions between PMC and PML guarantees and their relationships with other definitions, specifically ALIP, LIP, and LDP. We begin with finite random variables in Section V-A and demonstrate that if a mechanism satisfies a PML guarantee with a sufficiently small parameter, then it also satisfies a PMC guarantee. Conversely, we show that a PMC guarantee always implies a PML guarantee (Theorem 4). This result is then leveraged to establish connections between PML, PMC, LIP, ALIP, and LDP and derive bounds between these measures (Corollary 3). Here, we also highlight the practical implications of our results for designing mechanisms with optimal utility.
Next, in Section V-B, we extend our discussions to random variables on standard Borel spaces. First, we show that PMC is fundamentally more restrictive than PML in the sense that bounded PMC implies bounded PML (Theorem 5). We then prove an even stronger result stating that bounded PMC is equivalent to some form of an LDP guarantee (Theorem 6). This result demonstrates that PMC effectively serves as a context-aware [^1] A privacy measure is said to be context-aware if it depends on both the privacy mechanism and the distribution of the sensitive data. Conversely, it is said to be context-free if it depends on the privacy mechanism alone. version of LDP and its application is constrained to mechanisms compatible with LDP and DP. Overall, our findings in this section provide a clear and unified understanding of the relationships among privacy measures that bound the information density.
@@ I-C Other Related Works
The broader area of information theory has also contributed significantly to the study of privacy measures. Here, mutual information is notably the most extensively examined privacy measure [34, 35, 36, 37, 38, 39, 40]. Beyond mutual information, generalized information measures such as $f$-information [41] have also been used to assess privacy. Specific instances of $f$-information utilized as privacy measures include $\chi^{2}$-information [42, 43] (associated with $\chi^{2}$-divergence), and total variation privacy [44] (associated with total variation distance). Additionally, some works have used per-letter $f$-divergences (as opposed to the average-case divergence in $f$-information) as privacy measures [45, 46, 47]. Other research has examined the probability of correctly guessing [48] or guesswork [49, 50] for guaranteeing privacy.
Finally, several works have studied the design of privacy mechanisms subject to various privacy guarantees. These include local differential privacy [51, 52, 53], maximal leakage [54, 55], total variation privacy [44], and LIP [56, 57, 18], among others.
@@ II-A Notation and Assumptions
The following notations are used throughout the paper: $\mathbb{R}_{+}=[0,\infty)$ and ¯$\bar{\mathbb{R}}_{+}=[0,\infty]$ denote the non-negative real numbers and the extended non-negative real numbers. The set $[n]\coloneqq\{1,\ldots,n\}$ represents the first $n$ positive integers, $\log(\cdot)$ refers to the natural logarithm, and $\mathbf{1}_{\mathcal{A}}$ denotes the indicator function of the set $\mathcal{A}$. We also adopt the convention that $\nicefrac{{0}}{{0}}=1$.
@@
Let $\mathcal{X}$ be a set and $\mathcal{S}_{\mathcal{X}}$ a $\sigma$-algebra on $\mathcal{X}$. A mapping $X:\Omega\to\mathcal{X}$ is called a random variable taking values in $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ if $X$ is measurable relative to $\mathcal{H}$ and $\mathcal{S}_{\mathcal{X}}$. In this paper, we exclusively use $X$ to denote some data containing sensitive information, i.e., the secret. We use $P_{X}$ to denote the distribution of $X$. Often, we assume that $\mathcal{X}$ is a discrete set and use some notations specific to this case. In particular, (with a slight abuse of notation) we also use $P_{X}$ to denote the probability mass function (pmf) of $X$ and write $P_{X}(x)\coloneqq P_{X}(\{x\})$ for $x\in\mathcal{X}$. Moreover, we assume that $P_{X}$ has full support on $\mathcal{X}$, that is, $P_{X}(x)>0$ for all $x\in\mathcal{X}$, and define $p_{\mathrm{min}}\coloneqq\min\limits_{x\in\mathcal{X}}P_{X}(x)$.
@@
$\mathbb{E}[h]=\int_{\mathcal{X}\times\mathcal{Y}}h(x,y)\;P_{XY}(dx,dy)=\int_{\mathcal{X}}P_{X}(dx)\int_{\mathcal{Y}}h(x,y)\;P_{Y\mid X=x}(dy),$
@@
Finally, all measurable spaces are assumed to be standard Borel [58, Def. 8.35]. Standard Borel spaces have several convenient properties. Most prominently, joint distributions on standard Borel spaces can always be disintegrated into a kernel and a marginal distribution [59, Thm. IV.2.18]. That is, if $P_{XY}$ is a distribution on $(\mathcal{X}\times\mathcal{Y},\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{Y}})$ with marginal $P_{X}$ on $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$, then there exists a transition probability kernel from $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ into $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$ such that $P_{XY}=P_{X}\times P_{Y\mid X}$. Another advantage is that if $P_{Y\mid X}$ and $Q_{Y\mid X}$ are both kernels and $P_{Y\mid X=x}\ll Q_{Y\mid X=x}$ for all $x\in\mathcal{X}$, then we may invoke Doob’s version of the Radon-Nikodym theorem to obtain a Radon-Nikodym derivative $\frac{dP_{Y\mid X}}{dQ_{Y\mid X}}$ which is jointly measurable in $(x,y)$ [59, Thm. V.4.44]. This has the effect that we may alternatively use $\log\frac{dP_{Y\mid X}}{dP_{Y}}$ or $\log\frac{dP_{X\mid Y}}{dP_{X}}$ as the information density. Note that all measurable spaces of practical interest are standard Borel, e.g., countable sets, $\mathbb{R}^{n}$, or complete separable metric spaces endowed with Borel $\sigma$-algebras.
@@ II-B Pointwise Maximal Leakage
Pointwise maximal leakage (PML) is a recently introduced privacy measure that is operationally meaningful, robust, and flexible. For secrets with finite alphabets, PML is defined by quantifying risks in two powerful adversarial threat models: the gain function model, first introduced by Alvim et al. [14], and the randomized function model, put forward by Issa et al. [22]. It has been shown that when $X$ is a finite random variable, these two models are equivalent [12, Thm. 2]. Furthermore, the gain function model can be extended to secrets taking values in standard Borel spaces [13].
@@
On the other hand, if $\centernot\mathbb{P}\centernot{\ll}\mathbb{Q}$, then $D_{\infty}(\mathbb{P}\|\mathbb{Q})=\infty$. When the sample space $\Omega$ is countable we may write (4) in the form
@@
Definition 2 (Gain function view of PML [13, Def. 3]).
@@
$\ell_{P_{XY}}(X\to y)\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ g\in\mathcal{G}\end{subarray}}\frac{\sup\limits_{P_{W\mid Y}}\mathbb{E}\left[g(X,W)\mid Y=y\right]}{\sup_{w\in\mathcal{W}}\mathbb{E}[g(X,w)]},$ (5)
where the supremum in the numerator is over all transition probability kernels $P_{W\mid Y}$ from $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$ into $(\mathcal{W},\mathcal{S}_{\mathcal{W}})$, and $\mathcal{G}$ denotes the set of all gain functions defined as
$\mathcal{G}\coloneqq\left\{g\in(\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{W}})_{+}\Big{|}\sup_{w\in\mathcal{W}}\mathbb{E}[g(X,w)]<\infty\right\}.$
Definition 2 can be explained in the following way. Consider an adversary who aims to construct an estimate of $X$, denoted by $W$, in order to maximize a non-negative gain function $g$. The gain function $g$ encapsulates the adversary’s goal and can be adjusted to represent a variety of privacy attacks. For example, if $X$ is a dataset, then $g$ might model membership inference attacks [12]. To measure the information leakage associated with a single released outcome $y\in\mathcal{Y}$, we compare the expected value of $g$ after observing $y$ (in the numerator of (5)) with the expected value of $g$ before observing $y$ (in the denominator). The posterior expected gain is calculated using the optimal estimation kernel $P_{W\mid Y}$, while the prior expected gain is $\sup_{w\in\mathcal{W}}\mathbb{E}\left[g(X,w)\right]$. Then, to ensure that the privacy measure resulting from this definition encompasses various attacks, the ratio of the posterior-to-prior expected gain is maximized over all possible measurable spaces $(\mathcal{W},\mathcal{S}_{\mathcal{W}})$ and all gain functions in $\mathcal{G}$.
@@
Definition 3 (Randomized function view of PML [12, Def. 1]).
@@
$\ell_{P_{XY}}(X\to y)$ $=\log\sup_{U:U-X-Y}\,\frac{\sup_{P_{\hat{U}\mid Y}}\mathbb{P}\left[U=\hat{U}\mid Y=y\right]}{\max_{u\in\mathcal{U}}P_{U}(u)},$ (6)
where $U$ and $\hat{U}$ are random variables on a finite set $\mathcal{U}$, $P_{U}=P_{U\mid X}\circ P_{X}$, and the Markov chain $U-X-Y-\hat{U}$ holds.
Definition 3 can be understood similarly to Definition 2. Let $U$ be a randomized function of $X$. To quantify the information leakage associated with a single outcome of the mechanism $y\in\mathcal{Y}$, we compare the probability of correctly guessing $U$ after observing $y$ with the a priori probability of correctly guessing $U$ in a ratio. Then, to obtain a robust privacy measure, this posterior-to-prior ratio is maximized over all possible randomized functions of $X$, represented by the supremum over $U$’s satisfying the Markov chain $U-X-Y$.

### v2

@@ Abstract
This paper explores the implications of guaranteeing privacy by imposing a lower bound on the information density between the private and the public data. We introduce a novel and operationally meaningful privacy measure called pointwise maximal cost (PMC) and demonstrate that imposing an upper bound on PMC is equivalent to enforcing a lower bound on the information density. PMC quantifies the information leakage about a secret to adversaries who aim to minimize non-negative cost functions after observing the outcome of a privacy mechanism. When restricted to finite alphabets, PMC can equivalently be defined as the information leakage to adversaries aiming to minimize the probability of incorrectly guessing randomized functions of the secret. We study the properties of PMC and apply it to standard privacy mechanisms to demonstrate its practical relevance. Through a detailed examination, we connect PMC with other privacy measures that impose upper or lower bounds on the information density. These are pointwise maximal leakage (PML), local differential privacy (LDP), and (asymmetric) local information privacy. In particular, we show that a mechanism satisfies LDP if and only if it has both bounded PMC and bounded PML. Overall, our work fills a conceptual and operational gap in the taxonomy of privacy measures, bridges existing disconnects between different frameworks, and offers insights for selecting a suitable notion of privacy in a given application.
@@ I-A Motivations
There are two key motivations behind this paper. First, we seek to provide an in-depth understanding of the implications of defining privacy through a lower bound on the information density. To this end, we introduce an operationally meaningful privacy measure, which we call pointwise maximal cost (PMC). We demonstrate that imposing an upper bound on PMC is equivalent to imposing a lower bound on the information density. PMC may be thought of as a complementary definition to PML. More precisely, while PML considers opportunistic adversaries aiming to maximize gain functions, PMC focuses on risk-averse adversaries aiming to minimize arbitrary non-negative cost functions. The following simple example illustrates how the threat associated with opportunistic and risk-averse adversaries is quantified differently in a membership inference scenario [20].
Example 1.
Consider a membership inference scenario, where $X$ denotes a database and the adversary aims to determine whether or not a particular individual, Alice, is present in $X$. Let $\mathcal{X}$ denote the set of all possible databases, $\mathcal{X}_{1}\subset\mathcal{X}$ denote the set of databases that include Alice’s data, and $\mathcal{X}_{0}=\mathcal{X}\setminus\mathcal{X}_{1}$ denote the set of databases that do not include Alice’s data. Thus, each database $x\in\mathcal{X}$ satisfies $x\in\mathcal{X}_{i}$ for some $i\in\{0,1\}$, indicating Alice’s membership status.
In this context, an opportunistic adversary may be modeled via the gain function $g(w,x)=\mathbf{1}[w=i]$ for $x\in\mathcal{X}_{i}$, where $w\in\{0,1\}$ is the adversary’s guess, and $\mathbf{1}[\cdot]$ denotes the indicator function. Here, the adversary receives a gain of $1$ for correctly guessing the membership and no gain otherwise. The adversary would then seek a strategy for producing guesses that maximizes this gain.
On the other hand, a risk-averse adversary for the membership inference attack may be modeled via a cost function $c(w,x)=\mathbf{1}[w\neq i]$ for $x\in\mathcal{X}_{i}$ and $w\in\{0,1\}$. This adversary would seek a strategy for producing guesses that minimizes the cost, reflecting a preference to avoid incorrect guesses.
@@ I-B Contributions and Summary of the Main Results
Next, we argue that PMC can alternatively be defined as the multiplicative decrease in the expected value of an arbitrary non-negative cost function upon observing an outcome $y$ of the privacy mechanism. This setup is referred to as the cost function model. Essentially, we prove an equivalence between the randomized function model and the cost function model of PMC (Theorem 3). This equivalence signifies a parallelism with PML since a similar equivalence between the gain function model and the randomized function model of PML was previously established in [55, Thm. 2]. Table I provides a comparative summary of PMC and PML, with the relevant definitions presented side by side to highlight their similarities and differences.
TABLE I: Comparative summary of PMC and PML.
Pointwise Maximal Cost Pointwise Maximal Leakage
Quantifies the threat of risk-averse adversaries Definition via cost functions (Def. 11) $\Lambda_{P_{XY}}(X\to y)\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ c\in\mathcal{C}\end{subarray}}\;\frac{\inf\limits_{P_{W}}\mathbb{E}\Big[c(X,W)\Big]}{\inf\limits_{P_{\hat{W}\mid Y}}\mathbb{E}\Big[c(X,\hat{W})\mid Y=y\Big]}$ (2) Definition via randomized functions (Def. 7) $\Lambda_{P_{XY}}(X\to y)\coloneqq\sup_{U:U-X-Y}\log\frac{\inf\limits_{P_{\tilde{U}}}\;\mathbb{P}\Big[U\neq\tilde{U}\Big]}{\inf\limits_{P_{\hat{U}\mid Y}}\mathbb{P}\Big[U\neq\hat{U}\mid Y=y\Big]}$ When $\mathcal{X}$ is finite, cost and randomized function definitions are equivalent (Thm. 3). Computable form (Thm. 4) $\Lambda_{P_{XY}}(X\to y)=D_{\infty}(P_{X}\|P_{X\mid Y=y})$ Lower bounds the information density $\Lambda_{P_{XY}}(X\to y)\leq\varepsilon\iff-\varepsilon\leq i_{P_{XY}}(X;y)\;\;\text{a.s.}$ (3) $\Lambda_{P_{XY}}(X\to y)$ can be infinite, even when $\mathcal{X}$ is finite. Quantifies the threat of opportunistic adversaries Definition via gain functions [54, Def. 3] $\ell_{P_{XY}}(X\to y)\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ g\in\mathcal{G}\end{subarray}}\frac{\sup\limits_{P_{\hat{W}\mid Y}}\mathbb{E}\big[g(X,\hat{W})\mid Y=y\big]}{\sup\limits_{P_{W}}\mathbb{E}[g(X,W)]}$ (4) Definition via randomized functions [55, Def. 1] $\ell_{P_{XY}}(X\to y)=\sup_{U:U-X-Y}\log\frac{\sup\limits_{P_{\hat{U}\mid Y}}\mathbb{P}\left[U=\hat{U}\mid Y=y\right]}{\sup\limits_{P_{\tilde{U}}}\;\mathbb{P}\Big[U=\tilde{U}\Big]}$ When $\mathcal{X}$ is finite, gain and randomized function definitions are equivalent [55, Thm. 2]. Computable form [54, Thm. 3] $\ell_{P_{XY}}(X\to y)=D_{\infty}(P_{X\mid Y=y}\|P_{X})$ Upper bounds the information density $\ell_{P_{XY}}(X\to y)\leq\varepsilon\iff i_{P_{XY}}(X;y)\leq\varepsilon\;\;\text{a.s.}$ (5) $\ell_{P_{XY}}(X\to y)\leq\log\frac{1}{p_{\min}}$ when $\mathcal{X}$ is finite.
In Sections III-C and III-D, we explore the connections between PMC and other information measures designed with risk-averse adversaries in mind. These are maximal cost leakage [27, Def. 11], maximal realizable cost [27, Def. 12], and maximal guesswork leakage [36]. By carefully examining these definitions and their formulations, we observe that they are often more cumbersome and complex without strengthening the privacy model. For instance, both maximal cost leakage and maximal realizable cost involve a supremum over a Markov chain, which we demonstrate can be omitted. By contrast, we show that PMC offers a streamlined yet powerful and flexible formulation that effectively encompasses these alternative definitions. It is important to note that our discussions here are restricted to finite random variables, as the other measures have only been defined in this context.
@@
In Section V, we explore the interactions between PMC and PML guarantees and their relationships with other definitions, specifically ALIP, LIP, and LDP. We begin with finite random variables in Section V-A and demonstrate that if a mechanism satisfies a PML guarantee with a sufficiently small parameter, then it also satisfies a PMC guarantee. Conversely, we show that a PMC guarantee always implies a PML guarantee (Theorem 6). This result is then leveraged to establish connections between PML, PMC, LIP, ALIP, and LDP and derive bounds between these measures (Corollary 3).
We further show that a mechanism satisfies LDP if and only if it has bounded PMC for all priors with fixed support (Theorem 5). This result mirrors a known characterization of LDP in terms of PML (see Theorem 1) and also offers a novel interpretation of LDP through risk-averse adversaries. Table II summarizes key relationships between these privacy measures for finite random variables, including both results from prior work and new connections established in this paper.
Next, in Section V-B, we extend our analysis to random variables on standard Borel spaces. While in finite spaces a PMC guarantee always implies a PML guarantee, we show that this implication breaks down in more general settings. We construct two illustrative examples: one involving countably infinite random variables and another involving absolutely continuous ones. In both cases, the mechanisms have bounded PMC but unbounded PML. Lastly, we prove that a mechanism satisfies LDP if and only if it has both bounded PMC and bounded PML (Theorem 7). This shows that LDP is equivalent to simultaneously restricting the threat imposed by risk-averse and opportunistic adversaries. Overall, our findings in Section V provide a clear and unified understanding of the relationships among privacy measures that restrict the information density.
@@ I-C Other Related Works
The broader area of information theory has also contributed significantly to the study of privacy measures. Here, mutual information is notably the most extensively examined privacy measure [4, 6, 5, 63, 44, 40, 49]. Beyond mutual information, generalized information measures such as $f$-information [12] have also been used to assess privacy. Specific instances of $f$-information utilized as privacy measures include $\chi^{2}$-information [14, 62] (associated with $\chi^{2}$-divergence), and total variation privacy [48] (associated with total variation distance). Additionally, some works have used per-letter $f$-divergences (as opposed to the average-case divergence in $f$-information) as privacy measures [68, 67, 69]. Other research has examined the probability of correctly guessing [7] or guesswork [46, 45] for guaranteeing privacy. The design of privacy mechanisms under various privacy constraints has also received considerable attention. These include mechanisms for local differential privacy [33, 24, 41], maximal leakage [66, 53], total variation privacy [48], and local information privacy [26, 52, 30], among others.
Finally, we note that privacy is often one of several competing goals in reliable and trustworthy learning systems, alongside objectives such as good generalization capabilities [57], fairness [17], and adversarial robustness [43]. These goals have been jointly studied in various works. For instance, differential privacy has been linked to generalization in adaptive learning [16, 8]. Recent work has also explored the interplay between privacy and adversarial robustness [72].
TABLE II: Summary of key implications and equivalences between LDP, PML, PMC, LIP, and ALIP for finite alphabets.
Implications References
$\varepsilon$-LDP $\Rightarrow$ $\log\left(p_{\min}+e^{-\varepsilon}(1-p_{\min})\right)-$PML [30, Thm. 1]
$\varepsilon$-LDP $\Rightarrow$ $\log\left(p_{\min}+e^{\varepsilon}(1-p_{\min})\right)$-PMC Prop. 3
$\varepsilon$-PML with $\varepsilon<\log\frac{1}{1-p_{\min}}$ $\Rightarrow$ $\log\left(\frac{p_{\min}}{1-e^{\varepsilon}(1-p_{\min})}\right)$-PMC Thm. 6
$\varepsilon$-PMC $\Rightarrow$ $\log\left(\frac{1-e^{-\varepsilon}(1-p_{\min})}{p_{\min}}\right)$-PML Thm. 6
Equivalences References
$\varepsilon$-LIP $\Leftrightarrow$ $\varepsilon$-PMC and $\varepsilon$-PML Thm. 2, [55, Thm. 1]
$(\varepsilon_{1},\varepsilon_{2})$-ALIP $\Leftrightarrow$ $\varepsilon_{1}$-PMC and $\varepsilon_{2}$-PML Thm. 2, [55, Thm. 1]
$\varepsilon$-LDP $\Leftrightarrow$ $\varepsilon$-PML for all $P_{X}$ [27, Thm. 14], [22, Thm. 3]
$\varepsilon$-LDP $\Leftrightarrow$ $\varepsilon$-PMC for all $P_{X}$ Thm. 5
@@ II-A Notation and Assumptions
The following notations are used throughout the paper: $\mathbb{R}_{+}=[0,\infty)$ and $\bar{\mathbb{R}}_{+}=[0,\infty]$ denote the non-negative real numbers and the extended non-negative real numbers. The set $n]\coloneqq\{1,\ldots,n\}$ represents the first $n$ positive integers, $\log(\cdot)$ refers to the natural logarithm, and $\mathbf{1}_{\mathcal{A}}$ denotes the indicator function of the set $\mathcal{A}$. We also adopt the convention that $\nicefrac{{0}}{{0}}=1$.
@@
Let $\mathcal{X}$ be a set and $\mathcal{S}_{\mathcal{X}}$ a $\sigma$-algebra on $\mathcal{X}$. A mapping $X:\Omega\to\mathcal{X}$ is called a random variable taking values in $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ if $X$ is measurable relative to $\mathcal{H}$ and $\mathcal{S}_{\mathcal{X}}$. In this paper, we exclusively use $X$ to denote some data containing sensitive information, i.e., the secret. We use $P_{X}$ to denote the distribution of $X$. Often, we assume that $\mathcal{X}$ is a discrete set and use some notations specific to this case. In particular, (with a slight abuse of notation) we also use $P_{X}$ to denote the probability mass function (pmf) of $X$ and write $P_{X}(x)\coloneqq P_{X}(\{x\})$ for $x\in\mathcal{X}$. Moreover, we assume that $P_{X}$ has full support on $\mathcal{X}$, that is, $P_{X}(x)>0$ for all $x\in\mathcal{X}$, and define $p_{\mathrm{min}}\coloneqq\min\limits_{x\in\mathcal{X}}P_{X}(x)$ for finite $\mathcal{X}$.
@@
$\mathbb{E}[h]$ $=\int_{\mathcal{X}\times\mathcal{Y}}h(x,y)\;P_{XY}(dx,dy)$
$=\int_{\mathcal{X}}P_{X}(dx)\int_{\mathcal{Y}}h(x,y)\;P_{Y\mid X=x}(dy),$
@@
Finally, all measurable spaces are assumed to be standard Borel [35, Def. 8.35]. Standard Borel spaces have several convenient properties. Most prominently, joint distributions on standard Borel spaces can always be disintegrated into a kernel and a marginal distribution [10, Thm. IV.2.18]. That is, if $P_{XY}$ is a distribution on $(\mathcal{X}\times\mathcal{Y},\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{Y}})$ with marginal $P_{X}$ on $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$, then there exists a transition probability kernel from $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ into $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$ such that $P_{XY}=P_{X}\times P_{Y\mid X}$. Another advantage is that if $P_{Y\mid X}$ and $Q_{Y\mid X}$ are both kernels and $P_{Y\mid X=x}\ll Q_{Y\mid X=x}$ for all $x\in\mathcal{X}$, then we may invoke Doob’s version of the Radon-Nikodym theorem to obtain a Radon-Nikodym derivative $\frac{dP_{Y\mid X}}{dQ_{Y\mid X}}$ which is jointly measurable in $(x,y)$ [10, Thm. V.4.44]. This has the effect that we may alternatively use $\log\frac{dP_{Y\mid X}}{dP_{Y}}$ or $\log\frac{dP_{X\mid Y}}{dP_{X}}$ as the information density. Note that all measurable spaces of practical interest are standard Borel, e.g., countable sets, $\mathbb{R}^{n}$, or complete separable metric spaces endowed with Borel $\sigma$-algebras.
@@ II-B Pointwise Maximal Leakage
Pointwise maximal leakage (PML) is a recently introduced privacy measure that is operationally meaningful, robust, and flexible. For secrets with finite alphabets, PML is defined by quantifying risks in two powerful adversarial threat models: the gain function model, first introduced by Alvim et al. [3], and the randomized function model, put forward by Issa et al. [27]. It has been shown that when $X$ is a finite random variable, these two models are equivalent [55, Thm. 2]. Furthermore, the gain function model can be extended to secrets taking values in standard Borel spaces [54].
@@
[…]
@@
Definition 2 (Gain function view of PML [54, Def. 3]).
@@
$\ell_{P_{XY}}(X\to y)\coloneqq\log\sup_{\begin{subarray}{c}(\mathcal{W},\mathcal{S}_{\mathcal{W}}),\\ g\in\mathcal{G}\end{subarray}}\frac{\sup\limits_{P_{\hat{W}\mid Y}}\mathbb{E}\left[g(X,\hat{W})\mid Y=y\right]}{\sup\limits_{P_{W}}\,\mathbb{E}[g(X,W)]},$ (9)
where the supremum in the numerator is over all transition probability kernels $P_{\hat{W}\mid Y}$ from $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$ into $(\mathcal{W},\mathcal{S}_{\mathcal{W}})$, and $\mathcal{G}$ denotes the set of all gain functions defined as
$\mathcal{G}\coloneqq\left\{g\in(\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{W}})_{+}\Big|\sup_{P_{W}}\,\mathbb{E}[g(X,W)]<\infty\right\}.$
Definition 2 can be explained in the following way. Consider an adversary who aims to construct an estimate of $X$ in order to maximize a non-negative gain function $g$. The gain function $g$ encapsulates the adversary’s goal and can be adjusted to represent a variety of privacy attacks. For example, if $X$ is a dataset, then $g$ might model membership inference attacks [55]. To measure the information leakage associated with a single released outcome $y\in\mathcal{Y}$, we compare the expected value of $g$ after observing $y$ (in the numerator of (9)) with the expected value of $g$ before observing $y$ (in the denominator). The posterior expected gain is calculated using the optimal estimation kernel $P_{\hat{W}\mid Y}$, while the prior expected gain is $\sup_{P_{W}}\,\mathbb{E}[g(X,W)]$. Then, to ensure that the privacy measure resulting from this definition encompasses various attacks, the ratio of the posterior-to-prior expected gain is maximized over all possible measurable spaces $(\mathcal{W},\mathcal{S}_{\mathcal{W}})$ and all gain functions in $\mathcal{G}$.
@@
Definition 3 (Randomized function view of PML [55, Def. 1]).
@@
$\ell_{P_{XY}}(X\to y)$ $=\log\sup_{U:U-X-Y}\,\frac{\sup\limits_{P_{\hat{U}\mid Y}}\mathbb{P}\left[U=\hat{U}\mid Y=y\right]}{\sup\limits_{P_{\tilde{U}}}\;\mathbb{P}\Big[U=\tilde{U}\Big]}$, (10)
where $U$, $\hat{U}$ and $\tilde{U}$ are random variables on a finite set $\mathcal{U}$, $P_{U}=P_{U\mid X}\circ P_{X}$, and the Markov chain $U-X-Y-\hat{U}$ holds.
Definition 3 can be understood similarly to Definition 2. Let $U$ be a randomized function of $X$. To quantify the information leakage associated with a single outcome of the mechanism $y\in\mathcal{Y}$, we compare the probability of correctly guessing $U$ after observing $y$ with the a priori probability of correctly guessing $U$ in a ratio. Then, to obtain a robust privacy measure, this posterior-to-prior ratio is maximized over all possible randomized functions of $X$, represented by the supremum over $U$’s satisfying the Markov chain UXYU-X-Y.